OpenStudy (anonymous):

Find the coordinates of the midpoint of AB . A (4, 7) and B (-6, 6)? help please...

OpenStudy (anonymous):

Midpoint = \[\frac{ x(1)+x(2) }{ 2 },\frac{ y(1)+y(2) }{ 2}\] Therefore: \[\frac{ 4+(-6) }{ 2 },\frac{ 7+6 }{ 2 }\] Therefore: \[-1,\frac{ 13 }{ 2 }\] Is the midpoint of the line.

OpenStudy (anonymous):

thank you so much! is make sense! :)

OpenStudy (anonymous):

i need help another though, is okay?